OpenStudy (anonymous):

Quadratic formula: show work please!!! 1. 8y^2-22y-21

jimthompson5910 (jim_thompson5910):

\[\Large y = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large y = \frac{-(-22)\pm\sqrt{(-22)^2-4(8)(-21)}}{2(8)}\] \[\Large y = \frac{22\pm\sqrt{484-(-672)}}{16}\] \[\Large y = \frac{22\pm\sqrt{1156}}{16}\] \[\Large y = \frac{22+\sqrt{1156}}{16} \ \text{or} \ y = \frac{22-\sqrt{1156}}{16}\] \[\Large y = \frac{22+34}{16} \ \text{or} \ y = \frac{22-34}{16}\] \[\Large y = \frac{56}{16} \ \text{or} \ y = \frac{-12}{16}\] \[\Large y = \frac{7}{2} \ \text{or} \ y = -\frac{3}{4}\] So the solutions are \[\Large y = \frac{7}{2} \ \text{or} \ y = -\frac{3}{4}\]